# A Kähler and quaternion-Kähler spacetime structure

### 16 Mar 2026 v1

@@ 1. There are no half-integer spin elementary states for the extended Poincaré groups;
In the past, the fact that the pseudo-unitary algebras do not possess elementary spinor representations has led some authors to dismiss them as spectrum-generating algebras. However the non-existence of spinor states, as linear irreducible representations, is their most interesting feature. The nonexistence of spinors as linear representations, does not mean that they cannot be implemented in a non-linear way. This may be done by expressing the groups as principal bundles, with structure groups the subgroups of the real submanifolds, and then constructing the associated vector bundles. This has been carried out in [4] for massive and massless states. For example, for massive states the isotropy group is $U\left(3,\mathbb{C}\right)$ for the larger group and $SO\left(3,\mathbb{R}\right)$ for the real submanifold. $U\left(3,\mathbb{C}\right)$ is looked at as a principal bundle with base $W=U\left(3,\mathbb{C}\right)/SO\left(3,\mathbb{R}\right)$ and fiber $SO\left(3,\mathbb{R}\right)$. $W$ is not a spin manifold and to convert it into a spin manifold one has to add to it, by a Whitney sum, an additional $Sp\left(1\right)\simeq SU\left(2\right)$ manifold. In the process new quantum numbers are added to the spinors of the associated vector bundle. In addition one obtains a degeneracy of states with identical quantum numbers in the base $W$.
@@
In conclusion: requiring in this way the states in the real submanifold to inherit the symmetries of the larger complex group, a set of additional quantum numbers was obtained that is evocative of those of one-generation standard model. As stated in [4] the structure is not exactly the same as the algebra of the standard model. In particular the color-like degenerate states are not linear representations of $SU\left(3\right)$, but independent directions in the base manifold $W.$
@@
Is the quantum number structure of the standard model really a consequence of the symmetries of a larger space where our real spacetime is embedded? Is this an actual embedding or a mathematical fiction? Can it ever be decided? Strangely, so far the best hint comes from the algebraic structure of standard model.

### 31 Mar 2026 v2

@@ 1. There are no half-integer spin elementary states for the extended Poincaré groups;
In the past, the fact that the pseudo-unitary algebras do not possess elementary spinor representations has led some authors to dismiss them as spectrum-generating algebras. However the non-existence of spinor states, as linear irreducible representations, is their most interesting feature. The nonexistence of spinors as linear representations, does not mean that they cannot be implemented in a non-linear way. This may be done by expressing the groups as principal bundles, with structure groups the subgroups of the real submanifolds, and then constructing the associated vector bundles. This has been carried out in [4] [5] for massive and massless states. For example, for massive states the isotropy group is $U\left(3,\mathbb{C}\right)$ for the larger group and $SO\left(3,\mathbb{R}\right)$ for the real submanifold. $U\left(3,\mathbb{C}\right)$ is looked at as a principal bundle with base $W=U\left(3,\mathbb{C}\right)/SO\left(3,\mathbb{R}\right)$ and fiber $SO\left(3,\mathbb{R}\right)$. $W$ is not a spin manifold and to convert it into a spin manifold one has to add to it, by a Whitney sum, an additional $Sp\left(1\right)\simeq SU\left(2\right)$ manifold. In the process new quantum numbers are added to the spinors of the associated vector bundle. In addition one obtains a degeneracy of states with identical quantum numbers in the base $W$.
@@
In conclusion: requiring in this way the states in the real submanifold to inherit the symmetries of the larger complex group, a set of additional quantum numbers was obtained that is evocative of those of one-generation standard model. As stated in [4] [5] the structure is not exactly the same as the algebra of the standard model. In particular the color-like degenerate states are not linear representations of $SU\left(3\right)$, but independent directions in the base manifold $W.$
@@
Is the quantum number structure of the standard model really a consequence of the symmetries of a larger space where our real spacetime is embedded? Is this an actual embedding or a mathematical fiction? Can it ever be decided? Strangely, so far the best hint comes from the algebraic structure of standard model (see the Appendix).
@@ 3 The quaternion-Kähler structure
Appendix: Quaternionic spacetime and the standard model quantum numbers
According to the complex or quaternionic embeddings, massive color states exist as degrees of freedom in a (symmetric space) $6$ or 1818-dimensional manifolds ($U\left(3,\mathbb{C}\right)/SO\left(3\right)$ or $Sp\left(3\right)/SO\left(3\right)$), acted upon by $U\left(3,\mathbb{C}\right)$ or $Sp\left(3\right)$ gauge groups.
In the complex case, the base manifold for massive states $U\left(3,\mathbb{C}\right)/SO\left(3\right)$ the $6$ dimensions are compatible with one generation of colour quarks and anti-quarks and, in the quaternionic case $Sp\left(3\right)/SO\left(3\right)$ the $18$ dimensions allow for three generations of colour quarks and anti-quarks. Both $U\left(3,\mathbb{C}\right)$and $Sp\left(3\right)$ have no spinors as elementary representation spaces and if a spin $\frac{1}{2}$ state ($\psi_{1}$) is attached to the $SO\left(3\right)$ fibers, extra $SU\left(2\right)$ bundles (with the same base) should be added, as Whitney sums. A $SU\left(2\right)$ spinor ($\psi_{2}$) in the additional bundle, coupled with the $\psi_{1}$ spinor provides the spin $1$ of the $U\left(3\right)$ (and $Sp\left(3\right)$) representations, therefore realizing the matching of the symmetries of the real spacetime ($SO\left(1,3\right)$) with those of the embedding complex ($U\left(1,3,\mathbb{C}\right)$) or quaternionic ($U\left(1,3,\mathbb{H}\right)$) spacetimes. In addition, the extra $SU\left(2\right)$ bundles also endow the spinor states with a pair of flavour-like quantum numbers.
From the coupling of the two spinors a scalar field is also obtained. Hence one obtains two Higgs-like states for each generation. Notice that the requirement of compatibility of the symmetries of the real spacetime with those of the larger embedding spacetimes only requires the matching of the representations. In particular the additional $SU\left(2\right)$ bundles, might be different for each generation, leading to different mass structures.
For massless states the base manifolds would be $U\left(2\right)/SO\left(2\right)$ or $Sp\left(2\right)/SO\left(2\right)$ (complex or quaternionic embeddings) and the matching of the representations also requires Whitney sums generating flavour-like quantum numbers, but no colour degeneracy. In the quaternionic case the scalars that are obtained might again be distinct for each generation, without spoiling the symmetry of the embedding spacetime, generating different masses for different generations. The small masses of the neutrinos might have a different origin [26].